OpenStudy (anonymous):

I need to find the derivative of y=3x(x^3/2). I know the answer is 2x(3x^2). I just dont know how to get it. help please

OpenStudy (lalaly):

first multiply 3x and x^(3/2)

OpenStudy (anonymous):

My teacher said that you need to use the product rule. which is the first times the deriviative of the second, plus the second times the derivative of the first.

OpenStudy (anonymous):

Yup that is correct

OpenStudy (anonymous):

derivative of 3x=3

OpenStudy (anonymous):

Sorry the part isnt raised to 3/2 its x^3 divided by 2

OpenStudy (anonymous):

and X^(3/2)= (3/2)x^(1/2)

OpenStudy (anonymous):

YOu forgot to subtract one from the power

OpenStudy (lalaly):

im sorry, rld is right, i got distracted

OpenStudy (anonymous):

ya with the chat conversation!! :D

OpenStudy (lalaly):

lol :( sorry

OpenStudy (lalaly):

do u still need help with this, do u want me to do it again? no distractions:D

OpenStudy (anonymous):

3*(x^(2/3))+(2/3)x^(1/2)*3x 3x^(2/3)+2^(3/2)

OpenStudy (anonymous):

oh whoops io made a mistake

OpenStudy (anonymous):

please. i know how to do0 it i just think im making a simple mistake

OpenStudy (anonymous):

3x^(2/3)+2x(x^(1/2))

OpenStudy (anonymous):

wait give me a sec

OpenStudy (anonymous):

did u see where i said it wasnt raised to 2/3?

OpenStudy (anonymous):

oh i see i was trying to see how u got that!!!

OpenStudy (anonymous):

ok let me start from the begginning again

OpenStudy (lalaly):

\[3x(x^{\frac{3}{2}})\]\[\frac{d}{dx}(3x(x^{\frac{3}{2}})=3x^{\frac{3}{2}}+3x(x^\frac{1}{2}) \times \frac{3}{2}\]

OpenStudy (anonymous):

haha i didnt realize it until afterwards

OpenStudy (anonymous):

I think the equation is different that is what maclovin is claiming

OpenStudy (anonymous):

that it is 3x((x^3)/2)

OpenStudy (lalaly):

\[=3x^{\frac{3}{2}} +\frac{9}{2}x^{\frac{3}{2}}\]

OpenStudy (lalaly):

oh

OpenStudy (anonymous):